A block of mass m (initially at rest) is dropped from a height h landing onto a spring with spring/force constant k. Determine the maximum distance y that the spring will be compressed?


1
Expert's answer
2021-05-05T10:56:24-0400

Answer

By enery conservation

"mgh=\frac{1}{2}ky^2\\
y=\sqrt{\frac{2mgh}{k}}"